OpenStudy (anonymous):

Divide and simply to lowest terms. 4x^2-16x/ 4x-3 / x-4/ 4x-3

jhonyy9 (jhonyy9):

hope so much that you know sure how need dividing two factors

jhonyy9 (jhonyy9):

do you know it ?

jhonyy9 (jhonyy9):

so multiplie by second inversed ok ?

OpenStudy (anonymous):

no, which is why I am asking. From what my chapter states, a/b / c/d = a/b * d/c

jhonyy9 (jhonyy9):

why ? this is right sure a/b/c/d = a/b *d/c

OpenStudy (anonymous):

okay got that part, but now I have to simply 4x^2-16x/ x-4

jhonyy9 (jhonyy9):

yes so if you rewrite it what will get ?

OpenStudy (anonymous):

I am not sure, I know the LCM would be either x or 4

jhonyy9 (jhonyy9):

rewrite it with this property that you need multiplie the first fraction with second inversed and check what will get and what will can simplifie there

OpenStudy (anonymous):

so then it would be 4x(x-4)/ x-4 ?

jhonyy9 (jhonyy9):

yes and multiplie by second inversed

OpenStudy (anonymous):

so really it just equals 4x

jhonyy9 (jhonyy9):

4x(x-4) 4x -3 ------- * ------- = x-4 x-4

OpenStudy (anonymous):

4x(x-4) * 4x-3

jhonyy9 (jhonyy9):

not right

jhonyy9 (jhonyy9):

where is the denominator ?

OpenStudy (anonymous):

4x(x-4)*4x-3 over x-4

jhonyy9 (jhonyy9):

not right

jhonyy9 (jhonyy9):

you have simpolified from numerator the x-4 yes ?

OpenStudy (anonymous):

okay

jhonyy9 (jhonyy9):

so than what will get ?

OpenStudy (anonymous):

4x?

jhonyy9 (jhonyy9):

4x(x -4) 4x-3 4x(4x-3) -------- * ------ = --------- = 4x 4x-3 x -4 4x -3 yes you are right good luck bye

OpenStudy (anonymous):

thank you.

jhonyy9 (jhonyy9):

was my pleasure hope so much that is understandably sure right bye - bye